A local hamburger shop sold a combined total of 759 hamburgers and cheeseburgers on Wednesday. There were 59 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Wednesday?
Answer by nerdybill(7384)

A local hamburger shop sold a combined total of 759 hamburgers and cheeseburgers on Wednesday. There were 59 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Wednesday?
.
Let x = number of hamburgers sold
then
x+59 = number of cheeseburgers sold
.
x + x+59 = 759
2x+59 = 759
2x = 700
x = 350 (hamburgers sold)